OpenStudy (anonymous):

A toy company took a sample of its toys and found that 3% of the toys were defective. If the company shipped 1200 toys, how many of the toys would the company expect to be defective?

OpenStudy (anonymous):

would it be 120 or 12 perhaps?

OpenStudy (anonymous):

Do proprtions\[\frac{ 3percent }{ 100percent }=\frac{ xtoys }{ 1200toys }=>100x=3600=>x=\frac{ 3600 }{ 100 }=?\]